# Lower Bounds for Chain-of-Thought Reasoning in Hard-Attention Transformers

### v2

@@ Abstract
Chain-of-thought reasoning and scratchpads have emerged as critical tools for enhancing the computational capabilities of transformers. While theoretical results show that polynomial-length scratchpads can extend transformers’ expressivity from $TC^{0}$ to $PTIME$, their required length remains poorly understood. Empirical evidence even suggests that transformers need scratchpads even for many problems in $TC^{0}$, such as Parity or Multiplication, challenging optimistic bounds derived from circuit complexity. In this work, we initiate the study of systematic lower bounds for the number of CoT steps across different algorithmic problems, in the hard-attention regime. We study a variety of algorithmic problems, and provide bounds that are tight up to logarithmic factors. Overall, these results contribute to emerging understanding of the power and limitations of chain-of-thought reasoning.
@@ 1 Introduction
We study this question in the unique-hard attention (UHAT) regime, a popular theoretical abstraction of self-attention, in which every attention head attends to the unique position where attention weights are maximized (e.g. Hahn, 2020; Hao et al., 2022; Barceló et al., 2024; Svete & Cotterell, 2024; Bergsträßer et al., 2024; Yang et al., 2024; Barceló et al., 2025). In this model, we rigorously prove that the CoT for Parity described above is optimal up to constant factors (Theorem 4.2). Besides Parity, we consider three other tasks (Multiplication, Median, Reachability) that are also empirically challenging for transformers to solve without CoT. Across tasks, we show that CoT lengths need to scale at least linearly in the size of the original problem. We further show that our bounds are optimal up to logarithmic factors. Overall, these lower bounds place broad constraints on the inference-time compute needed to enable transformers to solve these problems.
@@ 2.2 Model of Transformers
The UHAT model can be viewed as the limit of the Softmax model when one attention score $a_{i,s}$ far exceeds the others. If more than one $s$ attains the maximal attention score, ties are broken according to some fixed rule (e.g., choosing the left- or right-most match, Yang et al. (2024)). The output of the attention block is computed by weighting according to attention weights $\hat{a}_{i,j}^{(k,h)}$ ($j\leq i$) [^1] We assume causal masking in line with standard language model architectures, in order to allow autoregressive generation of the CoT. Our techniques are also applicable to the setting where the input itself is processed with bidirectional attention as assumed by Abbé et al. (2024) (Remark B.7)., and applying a linear transformation $V$ (“value”); these are then aggregated across heads and combined with a skip-connection:
@@ 2.3 Formalizing CoTs
We note that $\Xi$ is not restricted to be finite. Some theoretical work has allowed the CoT voabulary to grow with the input length (Bhattamishra et al., 2020b; Abbé et al., 2024); our lower-bounds are robust to this.
@@ 3 Results: Generic CoT Bounds
We show that the existence of a sublinear-length CoT has strong implications about the function computed.
@@ 4.1 Lower Bounds for Regular Languages
We trained transformers on Parity on lengths between 1 and 500 (Figure 2 and Appendix C.1.2; results averaged across three runs). We considered the full CoT from Figure 1 (dots along diagonal line). We also considered CoTs where only every $k$-th step of the CoT was provided, shortening the CoT by $1/k$ (dots below diagonal line). Across input lengths, the model is successful when $k\lessapprox 3$. We also verified that two LLMs (DeepSeek-R1, Guo et al. (2025), and o1-mini, Jaech et al. (2024)) produce CoTs of at least linear length (Appendix D).
@@ 4.2 Multiplication
We trained small transformers (2 layers, 2 heads) to multiply binary numbers with up to 16 digits, either directly, or with the $\mathcal{O}(N\log N)$ CoT from Theorem 4.6. Multiplication failed to generalize when no CoT was given (Figure 12), with difficulty driven by the high-sensitivity middle digits (Figure 3). The CoT succeeded at all lengths (Table 1). See Appendix C.2.4 for details. We also verified that two LLMs (DeepSeek-R1 (Guo et al., 2025) and o1-mini (Jaech et al., 2024)) produce CoTs of at least linear length (Appendix D).
@@ 5.2 Related Work
Our results concerning Parity are distinct from results on Subset Parities. A long string of results has established difficulty of learning subset parities – that is, functions $\chi_{U}$ for $U\subseteq\{1,\dots,N\}$ – in various setups, and established benefits for providing intermediate steps (e.g. Wies et al., 2022; Abbe et al., 2023; Kim & Suzuki, 2024; Anonymous, 2024). Intuitively, the difficulty of learning subset parities emerges from the fact that one has to identify the set $U$ from the exponentially-large power-set. In learning such functions, there is a provable benefit to providing intermediate steps in training (Wies et al., 2022; Abbe et al., 2023; Kim & Suzuki, 2024; Anonymous, 2024). However, this argument does not readily imply that causal transformers will find the Parity function difficult to learn and that CoTs will help there, because the Parity function applies to a very specific set ($U=\{1,\dots,N\}$). Our results thus are distinct from results on subset parities due to Wies et al. (2022); Abbe et al. (2023); Kim & Suzuki (2024); Anonymous (2024). Interestingly, Abbe et al. (2024) show that Parity may be easy to learn for MLPs with Rademacher initialization; however, this requires a weight norm substantially growing with $N$, making transformer length generalization unlikely (Huang et al., 2024).
@@ Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. The authors thank Satwik Bhattamishra, Will Merrill, Sophie Hao, Anej Svete, Franz Novak, Ryan Cotterell, and Andy Yang for discussion, and Yash Sarrof and Entang Wang for feedback on the paper.
@@ Appendix B General Theoretical Results
This is a stronger statement of Lemma 4 of Hahn (2020). That lemma only stated that the final output ${\bm{y}}_{n}^{(L)}$ was preserved in $T^{\prime}$. However, by the way $T^{\prime}$ is obtained in that proof333“We can thus remove layer 0, convert layer-1 activations ${\bm{y}}_{j}^{(1)}$ into layer-0 activations ${\bm{y}}_{j}^{(0)}$, and obtain a $(c\cdot(2^{c}kH+1))$-transformer performing the same computation as before when $\rho^{(3)}$ is applied.” (quoted from Hahn, 2020). Here, “$\rho^{(3)}$” corresponds to the restriction $\rho^{\prime}$ from the statement of the lemma., the stronger statement for all activations follows. We also note that the original statement considered $\Sigma=\{0,1\}$, but the proof transfers to arbitrary finite alphabets $\Sigma$ without change, except that $T^{\prime}$ is now a $c\cdot(|\Sigma|^{c}kH+1)$-transformer, rather than a $c\cdot(2^{c}kH+1)$-transformer.
@@
Let $T$ be a UHAT transformer operating over a finite alphabet. Take any is $C\in(0,1)$. Then there is $c\in\mathbb{N}$, $k\in\mathbb{N}$ such that, for each $N>k$, there is $\rho$ such that
@@
A reader might wonder why, instead of introducing Lemma B.4, one cannot simply apply Theorem 1 of (Hahn, 2020) independently to each layer of a Transformer. The reason is that the latter approach would only ensure the existence of $L$ independent restrictions – one for each layer –whereas Lemma B.4 guarantees that a single restriction fixes each $y_{i}^{(l)}$.
@@ C.2.2 Remark: Autoregressive Multiplication
It is sufficient to show the lower bound on average sensitivity; the claim about UHAT follows from the inclusion of UHAT in $AC^{0}$ and Boppana’s upper bound on the average sensitivity of $AC^{0}$ functions (Boppana, 1997). We refer to O’Donnell (2014) to background on Fourier analysis of Boolean functions. We consider the Fourier transform of $F_{N}$ as a Boolean function on $\{-1,1\}^{3N-1}$:
$\hat{F}_{N}(X,Y,M_{1\dots N-1})=\sum_{A,B\subseteq[1,N];C\subseteq[1,N-1]}\lambda_{A,B,C}\cdot\chi_{A,B,C}(X,Y,M)$ (28)
for unique coefficients $\lambda_{A,B,C}\in[-1,1]$ and
$\chi_{A,B,C}(X,Y,M):=\prod_{i\in A}\hat{X}_{i}\prod_{j\in B}\hat{Y}_{j}\prod_{k\in C}\hat{M}_{k}$ (29)
We now investigate, for $d\in(0,1)$, the space $\mathfrak{F}_{d}$ of polynomials $\{-1,1\}^{3N-1}\rightarrow\{-1,1\}$ of degree at most $\leq 3dN$. Then
$\sup_{f\in\mathfrak{F}_{d}}\mathbb{E}_{X,Y}[f(X,Y,M(X,Y))\hat{F}_{N}(X,Y,M(X,Y))]$
$\leq$ $\sum_{A,B,C:|A|+|B|+|C|\leq 3dN}\left|\mathbb{E}_{X,Y}[\hat{F}_{N}(X,Y,M(X,Y))\chi_{A,B,C}(X,Y,M(X,Y))]\right|$
$\leq$ $2^{-DN}\sum_{l=1}^{3dN}{3N-1\choose 3dN}$
Choosing $d>0$ sufficiently small (in relation to $D$) makes this bound exponentially small in $N$. This entails that at least a constant fraction of the Fourier mass of $\hat{F}_{N}$ is on degrees of order $\sim N$. A linear average sensitivity follows. ∎
@@ C.3.2 Experiment
We use “;" to separate each number. As mentioned, we experiment with the full scratchpad and without any scratchpad, as well as every $k$th element in the full scratchpad, where $k=\{2,3,4,5,6,9,12\}$. The models are trained with cross entropy loss on tokens after SEP (excluding itself). The input length in figure 5 refers to number of tokens before and including SEP, and CoT length refers to number of token after SEP and before the final answer (last 3 digits). While models are trained to predict both the scratchpad and the answer, we do not consider scratchpad during testing. In other words, to evaluate the models, we give them tokens up to SEP, and let them generate with greedy search, and take the last generated number before EOS as the prediction. The prediction is considered correct only when it matches all digits in the answer. We run each experiment (each $N$ and each kind of scratchpad) with 3 random seeds, and report the average accuracy.
@@ Results.
Parity
Multiplication
Median
Figure 10: Length of reasoning traces of DeepSeek R1 (Guo et al., 2025) and o1-mini (Jaech et al., 2024) LLMs prompted to solve the tasks of parity (left), multiplication (center), and median (right). We only include traces leading to the correct answer. In both cases, the growth of CoT appears at least linear. Manual inspection of reasoning traces of DeepSeek R1 shows that the model perform the counting of all ones for Parity, and the naive quadratic algorithm for Multiplication.

### v3

@@ Abstract
Chain-of-thought reasoning and scratchpads have emerged as critical tools for enhancing the computational capabilities of transformers. While theoretical results show that polynomial-length scratchpads can extend transformers’ expressivity from $TC^{0}$ to $PTIME$, their required length remains poorly understood. Empirical evidence even suggests that transformers need scratchpads even for many problems in $TC^{0}$, such as Parity or Multiplication, challenging optimistic bounds derived from circuit complexity. In this work, we initiate the study of systematic lower bounds for the number of CoT steps across different algorithmic problems, in the hard-attention regime. We study a variety of algorithmic problems, and provide bounds that are tight up to logarithmic factors. Overall, these results contribute to emerging understanding of the power and limitations of chain-of-thought reasoning [^1] Code link: https://github.com/lacoco-lab/scratchpad_bounds.
@@ 1 Introduction
We study this question in the unique-hard attention (UHAT) regime, a popular theoretical abstraction of self-attention, in which every attention head attends to the unique position where attention weights are maximized (e.g. Hahn, 2020; Hao et al., 2022; Barceló et al., 2024; Svete & Cotterell, 2024; Bergsträßer et al., 2024; Yang et al., 2024; Barceló et al., 2025). Importantly, bounds in this regime entail bounds on realistic softmax transformers operating at fixed precision, as their expressiveness is upper-bounded by UHAT (Jerad et al., 2025).
In this model, we rigorously prove that the CoT for Parity described above is optimal up to constant factors (Theorem 4.2). Besides Parity, we consider three other tasks (Multiplication, Median, Reachability) that are also empirically challenging for transformers to solve without CoT. Across tasks, we show that CoT lengths need to scale at least linearly in the size of the original problem. We further show that our bounds are optimal up to logarithmic factors. Overall, these lower bounds place broad constraints on the inference-time compute needed to enable transformers to solve these problems.
@@ 2.2 Model of Transformers
The UHAT model can be viewed as the limit of the Softmax model when one attention score $a_{i,s}$ far exceeds the others. If more than one $s$ attains the maximal attention score, ties are broken according to some fixed rule (e.g., choosing the left- or right-most match, Yang et al. (2024); Jerad et al. (2025)). The output of the attention block is computed by weighting according to attention weights $\hat{a}_{i,j}^{(k,h)}$ ($j\leq i$) [^2] We assume causal masking in line with standard language model architectures, in order to allow autoregressive generation of the CoT. Our techniques are also applicable to the setting where the input itself is processed with bidirectional attention as assumed by Abbé et al. (2024) (Remark B.7)., and applying a linear transformation $V$ (“value”); these are then aggregated across heads and combined with a skip-connection:
@@ 2.3 Formalizing CoTs
We note that $\Xi$ is not restricted to be finite. Some theoretical work has allowed the CoT vocabulary to grow with the input length (Bhattamishra et al., 2020b; Abbé et al., 2024); our lower-bounds are robust to this.
@@ 3 Results: Generic CoT Bounds
We show that the existence of a sublinear-length CoT has strong implications for the function computed.
@@ 4.1 Lower Bounds for Regular Languages
We trained transformers on Parity on lengths between 1 and 500 (Figure 2 and Appendix C.1.2; results averaged across three runs). We considered the full CoT from Figure 1 (dots along diagonal line). We also considered CoTs where only every $k$-th step of the CoT was provided, shortening the CoT by $1/k$ (dots below diagonal line). Across input lengths, the model is successful when $k\lessapprox 3$. We also verified that two LLMs (DeepSeek-R1, Guo et al. (2025), and o1-mini, Jaech et al. (2024)) produce CoTs of at least linear length (Figure 6).
@@ 4.2 Multiplication
We trained small transformers (2 layers, 2 heads) to multiply binary numbers with up to 16 digits, either directly, or with the $\mathcal{O}(N\log N)$ CoT from Theorem 4.6. Multiplication failed to generalize when no CoT was given (Figure 12), with difficulty driven by the high-sensitivity middle digits (Figure 3). The CoT succeeded at all lengths up to 16 (Table 1). [^4] We focused extensive experiments to numbers of up this length due to resource constraints. See Appendix C.2.4 for details. We also verified that two LLMs (DeepSeek-R1 (Guo et al., 2025) and o1-mini (Jaech et al., 2024)) produce CoTs of at least linear length (Figure 6).
@@ 5 Discussion
Parity
Multiplication
Median
Figure 6: Length of reasoning traces of DeepSeek R1 (Guo et al., 2025) and o1-mini (Jaech et al., 2024) LLMs prompted to solve the tasks of parity (left), multiplication (center), and median (right). We only include traces leading to the correct answer. In both cases, the growth of CoT appears at least linear. Manual inspection of reasoning traces of DeepSeek R1 shows that the model perform the counting of all ones for Parity, and the naive quadratic algorithm for Multiplication. The experimental details are provided in Appendix D.
@@ 5.2 Related Work
Our results concerning Parity are distinct from results on Subset Parities. A long string of results has established difficulty of learning subset parities – that is, parity functions restricted to subsets $U\subseteq\{1,\dots,N\}$ – in various setups, and established benefits for providing intermediate steps (e.g. Wies et al., 2022; Abbe et al., 2023; Kim & Suzuki, 2024; Yang et al., 2025). Intuitively, the difficulty of learning subset parities emerges from the fact that one has to identify the set $U$ from the exponentially-large power-set. In learning such functions, there is a provable benefit to providing intermediate steps in training (Wies et al., 2022; Abbe et al., 2023; Kim & Suzuki, 2024; Yang et al., 2025). However, this argument does not readily imply that causal transformers will find the Parity function difficult to learn and that CoTs will help there, because the Parity function applies to a very specific set ($U=\{1,\dots,N\}$). Our results thus are distinct from results on subset parities due to Wies et al. (2022); Abbe et al. (2023); Kim & Suzuki (2024); Yang et al. (2025). Interestingly, Abbe et al. (2024) show that Parity may be easy to learn for MLPs with Rademacher initialization; however, this requires a weight norm substantially growing with $N$, making transformer length generalization unlikely (Huang et al., 2024).
Mechanistic Understanding of CoT
Concurrent with the theoretical line of work, the mechanisms behind the CoT abilities in LLMs have been studied empirically. Wang et al. (2024) used path patching to identify attention heads critical for CoT and found that only a few attention heads are important for that. Dutta et al. (2024) showed that LLMs employ multiple neural pathways that work in parallel to generate the final answer, all writing to the residual stream. Kudo et al. (2024) demonstrated for arithmetic reasoning that single-step subproblems can be solved by LLMs even before the scratchpad begins, while more complicated tasks are solved during CoT generation.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. The authors thank Satwik Bhattamishra, Will Merrill, Sophie Hao, Anej Svete, Franz Novak, Ryan Cotterell, Emmanuel Abbé, and Andy Yang for discussion, and Yash Sarrof and Entang Wang for feedback on the paper.
@@ Appendix B General Theoretical Results
This is a stronger statement of Lemma 4 of Hahn (2020). That lemma only stated that the final output ${\bm{y}}_{N}^{(L)}$ was preserved in $T^{\prime}$. However, by the way $T^{\prime}$ is obtained in that proof555“We can thus remove layer 0, convert layer-1 activations ${\bm{y}}_{j}^{(1)}$ into layer-0 activations ${\bm{y}}_{j}^{(0)}$, and obtain a $(c\cdot(2^{c}kH+1))$-transformer performing the same computation as before when $\rho^{(3)}$ is applied.” (quoted from Hahn, 2020). Here, “$\rho^{(3)}$” corresponds to the restriction $\rho^{\prime}$ from the statement of the lemma., the stronger statement for all activations follows. We also note that the original statement considered $\Sigma=\{0,1\}$, but the proof transfers to arbitrary finite alphabets $\Sigma$ without change, except that $T^{\prime}$ is now a $c\cdot(|\Sigma|^{c}kH+1)$-transformer, rather than a $c\cdot(2^{c}kH+1)$-transformer.
@@
Let $T$ be a UHAT transformer operating over a finite alphabet. Take any $C\in(0,1)$. Then there is $c\in\mathbb{N}$, $k\in\mathbb{N}$ such that, for each $N>k$, there is $\rho$ such that
@@
A reader might wonder why, instead of introducing Lemma B.4, one cannot simply apply Theorem 1 of Hahn (2020) independently to each layer of a Transformer. The reason is that the latter approach would only ensure the existence of $L$ independent restrictions – one for each layer –whereas Lemma B.4 guarantees that a single restriction fixes each $y_{i}^{(l)}$.
@@ C.2.2 Remark: Autoregressive Multiplication
It is sufficient to show the lower bound on average sensitivity; the claim about UHAT follows from the inclusion of UHAT in $AC^{0}$ and Boppana’s upper bound on the average sensitivity of $AC^{0}$ functions (Boppana, 1997). We refer to O’Donnell (2014) to background on Fourier analysis of Boolean functions. We consider the Fourier transform of $F_{N}$ as a Boolean function:
$\hat{F}_{N}(X,Y,M_{1\dots N-1})=\sum_{A,B\subseteq[1,N];C\subseteq[1,N-1]}\lambda_{A,B,C}\cdot\left[\prod_{i\in A}\hat{X}_{i}\prod_{j\in B}Y_{j}\prod_{k\in C}\hat{M}_{k}\right]$ (28)
for unique coefficients $\lambda_{A,B,C}\in\mathbb{R}$ given by
$\lambda_{A,B,C}=\mathbb{E}_{X,Y}\left[\hat{M}_{N}(X,Y)\prod_{i\in A}\hat{X}_{i}\prod_{j\in B}\hat{Y}_{j}\prod_{k\in C}\hat{M}_{k}(X,Y)\right]$ (29)
We have, by the conjecture,
$\lambda_{A,B,C}^{2}=$ $\left(\mathbb{E}\left[\hat{M}_{N}(X,Y)\prod_{i\in A}\hat{X}_{i}\prod_{j\in B}\hat{Y}_{j}\prod_{k\in C}\hat{M}_{k}(X,Y)\right]\right)^{2}\leq 2^{-2DN}$
Now, for any $\eta\in(0,1)$, the number of tuples $(A,B,C)$ where $|A|+|B|+|C|\leq\eta\cdot 3N$ is $\approx 2^{\gamma_{\eta}N}$ for some $\gamma_{\eta}>0$. Hence,
$\sum_{A,B,C:|A|+|B|+|C|\leq\eta\cdot 3N}|\lambda_{A,B,C}|^{2}\lessapprox 2^{(\gamma_{\eta}-2D)N}$
By Parseval’s Theorem:
$1=\sum_{A,B\subseteq[1,N];C\subseteq[1,N-1]}\lambda_{A,B,C}^{2}$
We choose $\eta$ so that $\gamma_{\eta}<2D$. Then, by the link between Fourier transforms and average sensitivity (O’Donnell, 2014), we find:
$as(F_{N})=$ $\sum_{A,B,C}\left(|A|+|B|+|C|\right)\cdot|\lambda_{A,B,C}|^{2}$
$\geq$ $3\eta N\sum_{A,B,C:|A|+|B|+|C|\geq\eta\cdot 3N}|\lambda_{A,B,C}|^{2}$
$\geq$ $3\eta N\cdot(1-2^{(\gamma_{\eta}-2D)N})$
$=$ $\Omega(N)$
This concludes the proof. ∎
@@ C.3.2 Experiment
We use “;” to separate each number. As mentioned, we experiment with the full scratchpad and without any scratchpad, as well as every $k$th element in the full scratchpad, where $k=\{2,3,4,5,6,9,12\}$. The models are trained with cross entropy loss on tokens after SEP (excluding itself). The input length in figure 5 refers to number of tokens before and including SEP, and CoT length refers to number of token after SEP and before the final answer (last 3 digits). While models are trained to predict both the scratchpad and the answer, we do not consider scratchpad during testing. In other words, to evaluate the models, we give them tokens up to SEP, and let them generate with greedy search, and take the last generated number before EOS as the prediction. The prediction is considered correct only when it matches all digits in the answer. We run each experiment (each $N$ and each kind of scratchpad) with 3 random seeds, and report the average accuracy.